# Symbiosis: Multi-Adapter Inference and Fine-Tuning

### v1

@@ 3.1 Design Goals
  2. 2.
    Flexible Placement: A client should be able to (a) share GPUs with the base executor (e.g., in resource constrained environment), (b) execute on a different GPU (to avoid interference), (c) execute on a CPU (e.g, to accommodate large KV cache) or (d) execute on another node (e.g., for privacy). Moreover, the client should transparently work with both sharded or non-sharded base model. This is particularly important to operate in an heterogeneous environment, consisting of different generations of GPUs. Symbiosis should be able to leverage such resources while minimizing their performance impact.
  3. 3.
@@
  4. 4.
    Batched Inference and Fine-tuning: Symbiosis should be batch inference and fine-tuning requests from different clients. This opens up an opportunity to improve GPU utilization.
  5. 5.
    Client Independence: The requests should be able to progress independently through model layers at different rates. This is because the clients with larger adapters (e.g., higher LoRA rank), longer sequence lengths, or located on slower accelerators require more time for an iteration, and need not be executed in lockstep with latency-sensitive requests.
  6. 6.
@@
vLLM [14] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
mLoRA [43] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Table 1: Design goals achieved in Symbiosis and other state-of-the-art multi-adapter systems. (FT: Fine-tuning)
@@
Figure 3: High-level design of Symbiosis
Figure 4: Symbiosis replaces the base model layers in the model structure on the client side with VirtLayer, which redirects their execution to the base executor.
Figure 5: The figure shows the various possible configurations of base executor and clients in Symbiosis.
3.2 Split Execution
Figure 3 shows the high-level architecture of Symbiosis. When loading the model, Symbiosis separates the base model layers as configured by the service provider and loads them to the base executor. Whereas the adapter and non-base layers, such as attention, batch normalization layer, are loaded by each tenant into respective client. The client and base executor can communicate through various mechanism shown in the diagram based on their host device. The service provider defines the batching policy for inference and fine-tuning requests. Tenant-specific privacy configuration is discussed below in Section 3.8.
@@ 3.2 Split Execution
A client can be a trainer for a fine-tuning job or an inference client for an inference job and serves as an endpoint to receive the training data or requests. Each client can opt for a preferred PEFT method, which is implemented through the adapter layers. Upon receiving input from a user, a client invokes the base executor for specific base model layers and sends it the activations for processing. Upon receiving a response from the base executor, it continues with the execution of the forward or backward pass for the remaining layers until the next base model layer. Eventually, the result is returned to the user (for inference), or an optimizer is invoked for model parameter update (for fine-tuning). Since each client drives its inference or training independently, the Symbiosis can accommodate different rates of execution for different adapters. For instance, client A can perform twice as many iterations as client B, where client A may share its execution of the base model layers with client B for a fraction of iterations.
To redirect the execution of the base model layers to the base executor, we scan and replace the frozen layers (base model layers) in the client-side model structure with a custom VirtLayer. Like all the layers in PyTorch, VirtLayer is an instance of torch.nn.Module with custom built forward and backward functions. These functions have the same properties, return datatypes, and sizes, as the corresponding functions of the replaced base model layer. When these functions are invoked, they send the necessary metadata, e.g., client id, target base layer, and the activation tensors to the base executor for processing. Upon receiving the response from the base executor, the function returns the received activations (forward pass) or gradients (backward pass) and the execution of the client-side layers continue. Since VirtLayer replaces the base model layer in the client-side model structure, this modification is performed through Symbiosis library call and the model code in the transformer library [35] need not change. Figure 4 shows the modified model structure.
@@
An invocation of an offloaded layer on the client-side results in an invocation of the base executor. The base executor serves each base model layer separately, so they can be invoked independently by different clients. The client-side passes the activations for the offloaded layer to the executor as tensors. The base executor batches the requests received from multiple clients and executes the requested layer’s forward or backward pass. In a forward pass of a fine-tuning job, the input and output tensors are saved to calculate the gradients in the corresponding backward pass. No tensors are saved for inference. Symbiosis introduces techniques (discussed in Section 3.6) to reduce, and in some cases eliminate, the storage required to store the input and output tensors.
3.3 Flexible Placement of Clients and Base Executor
@@
Symbiosis’s split execution decouples client and base executor resources, allowing for a flexible placement across different devices (or types of devices). Figure 5 shows the placement configurations supported by Symbiosis. The first configuration, local, show the case when one or more clients are located on the same GPU as the base executor. This configuration allows for fast communication while sharing the memory resources between clients and the base executor. In the remote configuration, clients are located on a different device than the base executor. The clients can be on another GPU on the same or a different node. It is suitable for clients that need higher GPU memory for large input batch sizes or sequence lengths. It also allows clients to scale independent of the base executor, where multiple GPUs can host several clients for the same base executor.
@@
Symbiosis also allows sharding models across GPUs. Sharding allows layers to be split across GPUs to reduce the memory footprint per GPU. Whenever a layer is executed in base executor, only the parameters corresponding to that layer are fetched from all the GPUs. After the layer’s execution, the fetched parameters are released, freeing the memory. The sharded local configuration allows scaling of the base model across multiple GPUs by sharding the base layer weights across them. A client can be present in any one of the GPUs where a shard of the base model resides. The base executor provides a communication endpoint at each of the GPU where its layers are sharded. Hence, a client only needs to communicate with its local shard. Lastly, the sharded remote configuration is similar to the remote configuration except that the base model layers are sharded across GPUs. This configuration allows executing the largest models, where both base executor and the clients can scale.
@@
To realize sharding, we utilize FSDP to shard the base model layers. FSDP enables sharding of parameters, gradient synchronization, and data-parallel training. Since all our base model layers are frozen (not trainable), we only use the sharding capabilities of FSDP. We design a custom wrapping strategy that marks individual base layers as independent FSDP instances. This enables base layers to scale and execute independent of the client layers, hence decoupling the execution of different layers.
3.4 Long-Context Inference with Heterogeneous Compute
@@
Symbiosis’s flexible placement and heterogeneous compute capabilities can be used to design an efficient system for long context inference. A major component of large inference runtime state in decoder models is the key-value (KV) cache, containing the key and value states used for calculating attention. The size of this KV cache can be huge (for example, 8GB for Llama2-7B model with an input sequence length of 16K and a batch size of just 1). Moreover, it grows linearly with input sequence length and batch size, often exceeding the memory available on a GPU. The state-of-the-art solutions offload the KV cache to CPU (exploiting the comparatively larger host memory), and saving GPU memory at the expense of increased CPU-GPU transfers [35]. However, for large output sequence lengths and/or batch sizes, the overhead of CPU-GPU transfers can become worse than the benefits of GPU acceleration as shown later in Figure 19.
@@
Figure 6 shows the execution of inference in Symbiosis. While we show client and base-executor GPUs separately for simplicity, any one of the placement configurations discussed earlier can be selected based on the model size and the number of clients. After the client layers and base model are loaded into the respective devices, we perform the throughput-sensitive task of prefill on GPU. The prefill stage processes the entire input prompt, requiring significant compute resources. We utilize the OffloadedCache proposed in [1] to implement a KV cache that is always offloaded to CPU. At any time, only the KV values corresponding to currently executing layers are cached on the GPU. During the computation of a base layer, we pre-fetch the corresponding KV values to the client-GPU, making them readily available for the attention calculation. Prefill is followed by the decoding stage, which generates one output token at a time. As a result, decoding is more latency sensitive, aiming to minimize mean token generation latency. To optimize for latency, we offload the client-side layers (and hence the compute) to the CPU during decoding. Since the KV cache and client layers are all present in the CPU memory, CPU performs all client-side decoding computations. We show in Section 4.3.2 that heterogeneous compute is faster than all-GPU compute for large outputs, owing to the high CPU-GPU transfer costs.
@@
Figure 6: Inference for large inputs in Symbiosis using flexible placement and heterogeneous compute. The wheels mark the devices performing computations during a stage.
3.5 Client - Base Executor Communication
@@
When client and base executor are located on the same GPU, Symbiosis implements a local communication mechanism. The clients and base executor communicate metadata and control messages over TCP via ZeroMQ [41], whereas the data is transferred through a shared tensor. Sharing obviates the need to transfer or copy the data and reduces communication latency. Sharing tensor across the processes is performed by acquiring the tensor metadata using share_memory_() method in the source process and rebuilding the reference using rebuild_cuda_tensor() method in the target process. However, the expensive CUDA calls for each layer can substantially increase the communication latency. To avoid this communication penalty, we pre-allocate a shared tensor for each client. This pre-allocated tensor is used for exchanging all input/output tensors between a server and a client.
@@
Figure 7: Per-layer wait time at the base executor for inference clients of Llama2-7B. In local configuration clients and base executor are on the same GPU, whereas for remote configuration the clients are on another GPU.
3.6 Independent Execution
PyTorch requires that the requests that were batched together for the forward pass of a layer be also be batched during their backward pass. This means that if inference and fine-tuning requests were batched together, both requests need to perform a backward pass. The inference requests do not require the backward pass and unnecessarily performing a backward pass would waste the GPU compute cycles. The limitation also means that if any fine-tuning requests that were batched together for a forward pass of a certain layer, should also be batched together for the backward pass of that layer. This limitation emerges from PyTorch’s implementation. When performing a forward pass, PyTorch captures the identity of the input/output tensors and the corresponding operation in a computation graph. This computation graph is traversed during the backward pass to compute gradients. PyTorch requires that the same tensors used in the forward pass to be used in the backward pass for gradient calculation. Therefore, the popular platforms, such as Transformers  [35], address this problem through lockstep execution, where a batch is maintained through the execution of forward and backward passes for all the layers.
@@ 3.6 Independent Execution
In Symbiosis, we eliminate the above batching requirement using the following two insights. First, since the layers on the base executor are frozen the parameters are not be updated during a backward pass. Second, for the most popular LLM architectures, (e.g., Llama, GPT, Bert), linear and 1D convolution layers constitute the largest fraction of the model parameters. For these layers, the input and output tensors are not involved in gradient calculations. Specifically, the gradient of output w.r.t. input resolves to the parameters themselves. Hence, Symbiosis does not store the input/output tensors for linear and 1D convolution base model layers, even for the fine-tuning request. Instead, it performs a matrix-multiplication between the output gradients and the parameters to generate the required gradients for these layers during the backward pass. This breaks the lockstep, while also providing significant memory savings by not requiring to store input/output tensors for each client at the base executor.
3.7 Opportunistic Batching
While breaking the lockstep execution frees each client to execute independently, this leads to smaller batches and requires the base executor to perform more iterations. To address this challenge, Symbiosisallows the base executor to wait to exploit the opportunity to accumulate new incoming requests at each layer and create larger batches. This is referred to as opportunistic batching. The wait time can be determined by the service provider based on the maximum tolerable penalty. To honor the latency of latency-sensitivity of inference request, we allow them to progress faster through the model without having to wait for less latency-sensitive requests for better batching.
@@ 3.7 Opportunistic Batching
3.8 Privacy for Multi-Tenancy
In multi-tenant environment, we consider a the following threat model. 1) The adversary’s goal is to extract the adapter parameters. The tenants that have trained their adapter on confidential data may not wish to expose their adapter parameters and request specific run-time state (KV cache) to the infrastructure provider or the base executor service provider. 2) Model extraction attack [34, 21, 12, 33], where the adversary steals the functionality of the model. With Symbiosis, it may be possible for the adversary to observe the communication from the client and base executor over a period of time and create an equivalent adapter.
We address both privacy concerns in the following ways. First, Symbiosis decouples the tenant specific state into a client process. Therefore, the client process can be hosted on a secure host deployed in tenant’s environment and it can communicate with the base model service over the network. Second, to address the privacy of activations, client adds noise to the activations before sending them to the base executor and the effect of noise is subtracted from the base executor’s output. This means the model produces the exact output which it otherwise would have in a non-privacy preserving setting. To accomplish this, we first send the noise to the base executor to calculate the noise effect. During regular execution, the base executor produces noisy output. The noise effect is subtracted from this noisy output to generate the actual output. The noise effect need not be calculated for every iteration, but only when once when for given noise value. To further prevent the leakage of noise, the tenant can either periodically change the noise or prepare several noise values in advance and pick different values for different iterations.
The calculate the effect of noise, we leverage the insight that common LLMs base layers are either $nn.Linear$ or $nn.Conv1D$ layers. For the layers that do not follow linearity, such as $Sigmoid$, $Softmax$, it is not possible to separate the effect of noise from the noisy output and produce the original output. Another challenge this privacy preserving technique is the presence of bias in $nn.Linear$ or $nn.Conv1D$ layers, which prevents the calculation of equation 2 below. To address this, we host another interface to the base executor which nullifies the effect of bias and returns $n_{effect}$.
Note that the privacy preservation itself does not address the security of the client environment, where users can rely on existing methods, such as NVIDIA MIG [2], to create a secure environment.
  1. 1.
    Calculation of noisy output.
    $y_{noisy}=Conv1D(x+n,W)+b$
    $y_{noisy}=Conv1D(x,W)+Conv1D(n,W)+b$
  2. 2.

### v2

@@ 3.1. Design Goals
  2. (2)
    Flexible Placement: A client should be able to (a) share GPUs with the base executor (e.g., in a resource constrained environment), (b) execute on a different GPU (to avoid interference), (c) execute on a CPU (e.g, to accommodate large KV cache) or (d) execute on another node (e.g., for privacy). This is particularly important to operate in an heterogeneous environment, consisting of different generations of GPUs. Symbiosis should be able to leverage such resources while minimizing their performance impact.
  3. (3)
@@
  4. (4)
    Batched Inference and Fine-tuning: Symbiosis should be able to batch inference and fine-tuning requests from different clients. This opens up an opportunity to improve GPU utilization.
  5. (5)
    Client Independence: The requests should be able to progress through the base model at different rates. This is because the clients with larger adapters (e.g., higher LoRA rank), longer sequence lengths, or located on slower GPUs or CPU require more time for an iteration, and need not be executed in lockstep with latency-sensitive requests.
  6. (6)
@@
vLLM (Kwon et al., 2023) $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
mLoRA (Zhengmao et al., 2023) $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Table 1. Design goals achieved in Symbiosis and other state-of-the-art multi-adapter systems. (FT: Fine-tuning)
@@
Figure 3. High-level design of Symbiosis
Figure 4. Symbiosis replaces the base model layers in the model structure on the client side with VirtLayer, which redirects their execution to the base executor.
Figure 5. The figure shows the various possible configurations of base executor and clients in Symbiosis. Even though Remote and Sharded Remote configurations show clients on the same GPU, they can also be placed on different GPUs or nodes.
3.2. Split Execution in Symbiosis
Design Overview
Figure 3 shows the high-level architecture of Symbiosis. When loading the model, Symbiosis separates the base model layers as configured by the service provider, and loads them to the base executor. Whereas the adapter and non-base layers, such as attention and batch normalization layers, are loaded by each tenant into respective client. The client and base executor can communicate through various mechanisms shown in Figure 3 based on their host devices. The service provider defines the batching policy for inference and fine-tuning requests. Tenants can enforce their resource management policies on the GPUs (or nodes) they have been allocated with the help of off-the-shelf schedulers (Fu et al., 2024; Wu et al., 2024a; Jain et al., 2025; Iliakopoulou et al., 2024). This allows them the flexibility to define per-client resource constraints (e.g., maximum context length, batch sizes) and reflect their priorities for different types of jobs. Section 3.8 discusses tenant-specific privacy configuration.
@@ 3.2. Split Execution in Symbiosis
A client can be a trainer (in case of a fine-tuning job) or an inference client (in case of an inference job) and serves as an endpoint to receive the training data or requests. Each client selects a PEFT method, with its desired parameters, which Symbiosis should implement through the adapter layers. This creates an instance of client-specific non-base layers, while the base model layers are served by the common base executor. For an input from a user, a client processes all client-side non-base layers locally and invokes the base executor for the base model layers. Whenever the client encounters a base model layer in the model, it sends the corresponding activations to the base executor for processing. Upon receiving a response from the base executor, the client continues with the execution of the client-side layers until it encounters the next base model layer. Both the forward and backward passes follow this layer-wise execution. Eventually, the result is returned to the user (for inference), or an optimizer is invoked for model parameter update (for fine-tuning). Since each client controls the rate at which it sends activations to the base executor, clients can drive their inference or training independently. This allows Symbiosis to accommodate different rates of execution for different adapters. For instance, client A can perform twice as many iterations as client B, where client A may share its execution of the base model layers with client B for a fraction of iterations.
@@
An invocation of a base model layer on the client-side results in an invocation of the base executor. The base executor serves each base model layer separately, so they can be invoked independently by different clients. The client-side passes the activations for the base model layer to the executor as tensors. The base executor batches the requests for a layer received from multiple clients and executes the requested layer’s forward or backward pass. The output of this batch processing is then split into individual outputs and sent to the respective clients. In a forward pass of a fine-tuning job, the input and output tensors are saved to calculate the gradients in the corresponding backward pass. No tensors are saved for inference. Symbiosis introduces techniques (discussed in Section 3.6) to reduce, and in some cases eliminate, the storage required to store the input and output tensors.
Symbiosis Virtlayer
To redirect the execution of the base model layers to the base executor, Symbiosis scans and replaces the frozen layers (base model layers) in the client-side model definition with a custom VirtLayer. Like all the layers in PyTorch, VirtLayer is an instance of torch.nn.Module with custom built forward and backward functions. These functions have the same properties, return datatypes, and sizes, as the corresponding functions of the replaced base model layer. However, these functions don’t execute the layers locally. Instead, when these functions are invoked, they send the necessary metadata, e.g., client id, target base layer, and the activation tensors to the base executor for processing. Upon receiving the response from the base executor, the functions return the received activations (forward pass) or gradients (backward pass) and the execution of the client-side layers continue. Each VirtLayer also has a base model layer identifier and other information needed to communicate with correct base model layer at the base executor. For example, VirtLayer contains client and base executor GPU identifiers (PyTorch ranks) in case of GPU-GPU communication. Since VirtLayer replaces the base model layer in the client-side model definition, this modification is performed through Symbiosis library call, eliminating the need to change the model code in the transformer library (Wolf et al., 2020a). Figure 4 shows the modified model definition.
3.3. Flexible Placement of Clients and Base Executor
@@
Symbiosis’s split execution decouples client and base executor resources, allowing for a flexible placement across different devices (or types of devices). Figure 5 shows the placement configurations supported by Symbiosis. The first configuration, local, show the case when one or more clients are located on the same GPU as the base executor. This configuration allows for fast communication while sharing the memory resources between clients and the base executor. In the remote configuration, clients are located on a different device than the base executor. The clients can be on another GPU on the same or a different node. It allows clients to scale independently of the base executor, where single or multiple GPUs can host several clients for the same base executor.
@@
Symbiosis also allows sharding models across GPUs. Sharding splits layers across GPUs to reduce the memory footprint per GPU. Whenever a layer is executed in base executor, only the parameters corresponding to that layer are fetched from all the GPUs. After the layer’s execution, the fetched parameters are released, freeing the memory. The sharded local configuration allows scaling of the base model across multiple GPUs by sharding the base layer weights across them. A client can be present in any one of the GPUs where a shard of the base model resides. The base executor provides a communication endpoint at each of the GPU where its layers are sharded. Hence, a client only needs to communicate with its local shard. Lastly, the sharded remote configuration is similar to the remote configuration except that the base model layers are sharded across GPUs. This configuration allows executing the largest models, where both base executor and the clients can scale independently.
@@
To realize sharding, we utilize Fully-Sharded Data Parallelism (FSDP) to shard the base model layers. FSDP enables sharding of parameters, gradient synchronization, and data-parallel training. Since all our base model layers are frozen (not trainable), we only use the sharding capabilities of FSDP. We design a custom FSDP wrapping strategy that marks individual base layers as independent FSDP instances. This enables base layers to scale and execute independent of the client layers, hence decoupling the execution of different layers.
3.4. Long-Context Inference with Heterogeneous Compute
@@
Symbiosis’s flexible placement and heterogeneous compute capabilities can be used to design an efficient system for long context inference. A significant fraction of inference runtime state in decoder models is the key-value (KV) cache, containing the key and value states used for calculating attention. The size of this KV cache can be large (for example, 8GB for Llama2-7B model with an input sequence length of 16K and a batch size of just 1). Moreover, it grows linearly with input sequence length and batch size, often exceeding the memory available on a GPU. The state-of-the-art solutions offload the KV cache to CPU (exploiting the comparatively larger host memory), and saving GPU memory at the expense of increased CPU-GPU transfers (Wolf et al., 2020a). However, for large output sequence lengths and/or batch sizes, the overhead of CPU-GPU transfers can become worse than the benefits of GPU acceleration as shown later in Figure 20.
@@
Figure 6 shows the execution of inference in Symbiosis. While we show client and base-executor are located on different GPUs, any of the placement configurations discussed earlier can be selected. After the client layers and base model are loaded into the respective devices, we perform the throughput-sensitive prefill on GPU. The prefill stage processes the entire input prompt, therefore it requires GPU acceleration to complete the prefill reasonably quickly. During prefill, we use the OffloadedCache feature proposed in  (off, 2024) to offload the KV cache to CPU memory. Prefill is followed by the decoding stage, which generates one output token at a time. Since the KV cache is already offloaded to the CPU memory, the client-side layers are also loaded and executed on the CPU. We show in Section 4.3.2 that such heterogeneous compute is faster than all-GPU compute, where KV cache needs to be transferred from CPU memory to GPU memory for generation. Moreover, since GPU cannot accommodate the entire KV cache, every iteration, the executing layer’s KV cache is fetched right before their execution.
@@
Figure 6. Inference for large inputs in Symbiosis using flexible placement and heterogeneous compute. The wheels mark the devices performing computations during a stage.
3.5. Client - Base Executor Communication
@@
When client and base executor are located on the same GPU, Symbiosis implements a local communication mechanism. The clients and base executor communicate metadata and control messages over TCP via ZeroMQ (ZeroMQ, 2024), whereas the data is transferred through a shared tensor. Sharing obviates the need to transfer or copy the data and reduces communication latency. Sharing tensor across the processes is performed by acquiring the tensor metadata using share_memory_() method in the source process and rebuilding the reference using rebuild_cuda_tensor() method in the target process. However, the expensive CUDA calls for each layer can substantially increase the communication latency. To avoid this communication penalty, we pre-allocate a shared tensor for each client. This pre-allocated tensor is used for exchanging all input/output tensors between a server and a client.
@@
Figure 7. Per-layer wait time at the base executor for inference clients of Llama2-7B. In local configuration clients and base executor are on the same GPU, whereas for remote configuration the clients are on another GPU.
3.6. Independent Execution
PyTorch requires that the requests that were batched together for the forward pass of a layer be also be batched during their backward pass. This means that if inference and fine-tuning requests were batched together, both requests need to perform a backward pass. The inference requests do not require the backward pass and unnecessarily performing a backward pass would waste the GPU compute cycles. The limitation also means that if any fine-tuning requests that were batched together for a forward pass of a certain layer, should also be batched together for the backward pass of that layer. This limitation emerges from PyTorch’s implementation. When performing a forward pass, PyTorch captures the identity of the input/output tensors and the corresponding operation in a computation graph. This computation graph is traversed during the backward pass to compute gradients. PyTorch requires that the same tensors used in the forward pass to be used in the backward pass for gradient calculation. Therefore, the popular platforms, such as Transformers  (Wolf et al., 2020a), address this problem through lockstep execution, where a batch is maintained through the execution of forward and backward passes for all the layers.
@@ 3.6. Independent Execution
In Symbiosis, we eliminate the above batching requirement using the following two insights. First, since the layers on the base executor are frozen the parameters are not be updated during a backward pass. Second, for the most popular LLM architectures, (e.g., Llama, GPT, Gemma, Bert), linear and 1D convolution layers constitute the largest fraction of the model parameters. For these layers, the input and output tensors are not involved in gradient calculations. Specifically, the gradient of output w.r.t. input resolves to the parameters themselves. Hence, Symbiosis does not store the input/output tensors for linear and 1D convolution base model layers, even for the fine-tuning request. Instead, it performs a matrix-multiplication between the output gradients and the parameters to generate the required gradients for these layers during the backward pass. This breaks the lockstep, while also providing significant memory savings by not requiring to store input/output tensors for each client at the base executor.
3.7. Opportunistic Batching
While breaking the lockstep execution frees each client to execute independently, this leads to smaller batches and requires the base executor to perform more iterations. To address this challenge, Symbiosis allows the base executor to wait to exploit the opportunity to accumulate new incoming requests and create larger batches. This is referred to as opportunistic batching. To honor the latency of latency-sensitivity of inference request, we allow them to progress faster through the model without having to wait for less latency-sensitive requests for better batching. To accomplish this, we base the wait time on the size of request. For instance, fine-tuning, prefill or a large batch of inference requests can afford to wait longer than smaller requests, since the wait time is a smaller fraction of their naturally longer iteration latency.
@@ 3.7. Opportunistic Batching
3.8. Privacy for Multi-Tenancy
Figure 8. In Symbiosis, the base executor has access to the activations marked in blue colored lines. Therefore, the effect of adapter parameters ($Wa$ $Wb$) can be computed with $(C-B)/A$.
In a multi-tenant environment, we consider the threat model where the adversary’s goal is to extract the adapter parameters. The tenants that have trained their adapter on confidential data may not wish to expose their adapter parameters and request-specific run-time state (KV cache) to the infrastructure provider or the base executor service provider. In Symbiosis, the third-party base executor can observe the client-side activations and steal its functionality, like the model extraction attacks (Wang et al., 2025; Liang et al., 2023; Gong et al., 2020; Tramèr et al., 2016). Figure 8 shows an example of the model architecture where the base executor can extract LoRA adapter parameters in Symbiosis.
We address the privacy concern in the following ways. First, Symbiosis decouples the tenant specific state into a client process. Therefore, the client process can be hosted on a secure host deployed in tenant’s environment and it can communicate with the base model service over the network. Second, to address the privacy of activations, client adds noise to the activations before sending them to the base executor and the effect of noise is subtracted from the base executor’s output. Note that due to subtraction, this mechanism does not alter the overall result, i.e., the model produces the exact output which it otherwise would have in a non-privacy preserving setting. To accomplish this, we first send the noise to the base executor to calculate the noise effect. During regular execution, the base executor produces noisy output. The noise effect is subtracted from this noisy output to generate the actual output. The noise effect need not be calculated for every iteration, but only once for a given noise value. To further prevent the leakage of noise, the tenant can either periodically change the noise or prepare several noise values in advance and pick different values for different iterations.
The calculate the effect of noise, we leverage the insight that common LLMs base layers are either $nn.Linear$ or $nn.Conv1D$ layers. For the layers that do not follow linearity, such as $Sigmoid$, $Softmax$, it is not possible to separate the effect of noise from the noisy output and produce the original output. Another challenge with this privacy preserving technique is the presence of bias in $nn.Linear$ or $nn.Conv1D$ layers, which prevents the calculation of equation 2 below. To address this, we also host an alternate execution flow in the base executor which nullifies the effect of bias and returns $n_{effect}$.
  1. (1)